Status Report

NASA Space Station On-Orbit Status 5 Apr 2004

By SpaceRef Editor
April 5, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 5 Apr 2004

All ISS systems continue to function nominally, except those noted previously or below.  Underway: Week 24 for Increment 8 (and 24 days until return of Expedition 8).

Update on Portable Computer System (PCS) Laptops:   After fruitless troubleshooting attempts over the weekend by the CDR, the shell of the PCS connected to the Cupola RWS (robotic workstation) was declared failed.  This leaves two working PCS machines on board, viz., one at the Lab RWS, the other in the Service Module (SM).  Applicable Flight Rule requirement is two active PSC minimum.  Current plans are to use crewtime on 4/7 (Wednesday) to prepare another PCS from a suitable shell of an SSC (station support computer) laptop.   [The failed Cupola PCS has shell #6060 and HD (hard drive) #6106.  Foale’s troubleshooting included changing out the HD (even trying an SSC HD) and the 1553 PCMCIA memory card, but early this morning (5:41am EDT), the PCS dropped off line again.  Turning an SSC shell into a PCS requires changing CMOS settings and HD swapout.  With appropriate testing this will likely take several hours but could replace an already planned PCS HD “ghosting” activity on 4/7.  A new ThinkPad 760XD is currently on low-priority standby for Progress 14P’s manifest.] 

Update on Noise:  TsUP/Moscow continues to look into the cause of the anomalous noise heard last week (4/2) in the vicinity of Foale’s sleep compartment in the SM.  Main suspect is a ventilation fan, and cleaning its screen will be scheduled ASAP.  However, TsUP may also have the general location of the noise inspected from the outside during EVA-10 (on 7/22).

Ku-band Antenna:  Evaluation of black marks seen on the outer edge of the Ku-band dish during the recent external station survey indicates the possibility that the antenna is contacting surrounding structure, perhaps (according to 3D modeling) a launch lock and/or a handrail.   [The Ku-band SGANT (space-to-ground antenna), a single beam boom with a dish, was deployed during the first EVA on 10/15/2000 on the Z1 integrated truss.  The mark was noticed first in April 2001.  Investigation is underway, and if there is indeed a clearance problem, a “mask” would be incorporated in the antenna steering program to exclude the interference region.]

FE Alexander Kaleri worked several hours on the SUBA onboard equipment control system behind SM wall panels, replacing three cable bundles with one cable trunk from FGB stowage.   [Five hours were set aside on his schedule for the activity, which included demating 12 connectors of the old jumpers, remating 12 connectors of the new cable trunk, re-labeling a KTK switch box (triple-command commutator), tying up cables, and covering the old connectors with protective covers.]

After ground-commanded activation of the MSG ESEM (Microgravity Science Glovebox/Exchangeable Standard Electronic Module), Mike Foale performed a visual inspection of the MSG and turned the facility on, prior to conducting the last of five scheduled PFMI (Pore Formation & Mobility Investigation) sessions, today on sample PFMI-5.   [After its start-up by Mike, untended operation of the experiment is underway for 11 hours.]

The Science Officer also conducted the fourth photography session of the BCAT-3 (Binary Colloidal Alloy Test-3) experiment, mounting the selected SGSM (Slow Growth Sample Module) in the MWA (Maintenance Work Area), photographing it sequentially with the Kodak 760 digital still camera and then stowing the sample.   [Because of the faster-than-expected homogenization of the alloy caused by the BCAT magnet’s field in micro-G, BCAT sessions were increased from originally planned three to five.]

Alex Kaleri worked with the CSA-CPs (compound specific analyzer-combustion products) for their monthly maintenance.  [He first conducted a calibration, to “re-zero” the gas sensors, then changed out the batteries of CSA-CPs #1009 & #1010.  The latter was designated as primary, the former as backup, both used for CO, HCl, and HCN only, while CSA-CP #1003 is available for measuring O2 only.]

Subsequently, the FE performed the periodic data take with both CSA-CPs.

Foale had time reserved for printing out recently uplinked CCS (command & control system) warning tables for updating two onboard Warning Books, and for making additional P&I (pen & ink) changes in the books for the CCS and the Lab IATS (internal active thermal control system).

The CDR’s planned activity of printing out an updated medical ACLS (advanced cardiac life support) algorithm was deferred.  [This algorithm is a flow-chart of procedural logic steps for aiding a crewmember in a cardio-vascular emergency, designed for two-man crews.]

Mike completed remaining IMAK (ISS medical accessory kit) transfers, relocating extra medical supplies, brought up in Soyuz 7S, to the proper medical kits.
The FE performed his daily hardware status check on the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.

The crew had their daily workouts on the TVIS treadmill, RED exerciser and VELO cycle with load trainer, and both crewmembers attended to the weekly TVIS maintenance by inspecting their individually set SPDs (subject positioning devices).

In addition, Mike performed the monthly treadmill maintenance as well as the regularly scheduled inspection and maintenance of the RED, including bolt tightening (done once a month).

The crew performed the daily routine maintenance of the SM’s SOZh life support system (including ASU toilet facilities), prepared the daily IMS “delta” file for updating the database, and conducted the regular routine status checkup of payloads currently running autonomously in the Lab (CGBA, PCG-STES010).

The CDR’s scheduled 18th weekly filling-out of the Food Frequency Questionnaire (FFQ) was deferred.   [The FFQ keeps a log of his nutritional intake over time on the medical equipment computer (MEC).]

At about 1:05pm, Mike had his weekly teleconference with the JSC Astronaut Office (Kent Rominger), via S-band S/G-2.

At 10:00am EDT, PCU-1 (plasma contactor unit #1), after being on Discharge for the past 11 hours, was shut down by ground command, then moded to nominal Standby.  PCU-2 will undergo the same shutdown and switchover to Standby tonight at 9:00pm.   [The PCU is the best currently available means to prevent charge buildup (“static electricity”) on spacecraft structures, which accumulates on them when high-voltage photovoltaic arrays (or ion thrusters, electrodynamic tethers, etc.) are operating in space.  The station’s two PCUs emit a stream of charged xenon atoms (ions) into space, representing a “ground” to the environment, which prevents large charges from building up on the structure.  They are activated for EVAs to prevent potential hazard of arcing damage or electrical shocks, although past measurements have indicated that spacewalkers are in no danger of electrical shocks and discharges.  The measurements have shown that the PCUs reduce the already low voltage by 3-5 volts.  Purpose of the temporary PCU activation for collecting data in XVV flight attitude was two-fold: (1) to provide the Plasma Hazard Evaluation Process (PHEP) Team with additional data needed to continue establishing a relationship between the charging hazard, ambient ionospheric conditions and the flight configuration/attitude; (2) the PCUs’ latch valve #2 has not been closed in over two years, since this normally happens only in the rarely entered Shutdown mode. On Flight 12A.1, however, the PCUs will be power-cycled during the EVA, requiring valve closure for recovery; thus, their operational status needed to be verified well before then.]

Looking ahead:   Starting Wednesday (4/7), each crewmember will have an additional hour per day for coming-home preparations.  Mike Foale will have five hours scheduled this week for clothing consolidation.  Also, this week the crew will be involved in slightly more than 12 hours of science/payload operations.

Today’s CEO (Crew Earth Observations) targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Logone Basin Fans, Chad (the complex of inland basins combined with long-term climate changes have produced numerous, poorly understood alluvial fan series in this region.  As ISS tracked across this target area, the crew was to try for a nadir mapping sequence to begin baselining the character and distribution of these features), Mexico City, Mexico (an excellent nadir pass over the sprawling Mexican capital), Atlanta, Georgia (weather should be clear by this pass.  The crew had a nadir view of this complex urban area), and Jarvis Island (400mm-lens.  A nadir pass in good weather and light over this isolated atoll.  Crew was to use the long lens for details of the coral reef structures).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of today, 1:59pm EDT).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair not completed; to be tested ASAP).  SFOG slot#2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.8; ppO2 (mmHg) — 160.4; ppCO2 (mmHg) — 3.7;
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.0.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 25.7.
  • Node:  Pressure (mmHg) — 758.52; temperature (deg C) — 24.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 760.07; temperature (deg C) — 25.1; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 760.17; temperature (deg C) — 24.8; shell heater temp (deg C) — 24.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — 25.0
  • PMA-2:  Shell heater temp (deg C) — 11.9

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, “sun slicer”, i.e., drag reduction-biased by 47 deg).
  • SM batteries:  Battery #8 is off-line; battery #6 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #6 is off-line (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Discharge mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3926 kg (8655 lb) as of 3/26/04;  [SM(755) + FGB(2512) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 4/14.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:48am EDT [= epoch]):

  • Mean altitude — 364.9 km
  • Apogee — 371.8km
  • Perigee — 358.0 km
  • Period — 91.84 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0010253
  • Solar Beta angle — -24.7 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.68
  • Mean altitude loss last 24 hours — 170 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 30689

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.